OpenStudy (anonymous):

Which of the following best describes the graph of f(x) = x2 - 3x - 10?

OpenStudy (anonymous):

Minimum at (1.5, -12.25) with intercepts at (5, 0) and (-2, 0) Minimum at (-1.5, -12.25) with intercepts at (-5, 0) and (2, 0) Minimum at (-3.5, -2.25) with intercepts at (-5, 0) and (-2, 0) Minimum at (3.5, -2.25) with intercepts at (5, 0) and (2, 0)

OpenStudy (ankit042):

you know how to proceed?

OpenStudy (anonymous):

no if i did i wouldnt be on here

OpenStudy (solomonzelman):

At first, @yeah_i_suck_at_math can you find the x intercepts of the given function ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

One thing that helps me a lot is the software called Geogebra. It's free and available at http://www.geogebra.org/cms/en/download/. You can just type a function like this into it to have it graphed. Can you tell us what an intercept is? What's an x-intercept? What's a y-intercept?

OpenStudy (solomonzelman):

@BillBell she should be able to find the x intercepts of this function without graphing.

OpenStudy (anonymous):

x axis y axis

OpenStudy (anonymous):

i should but i cant so you can leave if it bothers you

OpenStudy (solomonzelman):

Just by setting the function equal zero and solving for x.

OpenStudy (solomonzelman):

I'll start you off... f(x) = x^2 - 3x - 10 setting the function equal to zero, x^2 - 3x - 10 = 0 Can you solve for x now? (factoring would be good in this case)

OpenStudy (anonymous):

theres noway

OpenStudy (solomonzelman):

Can you factor x^2-3x-10=0 ?

OpenStudy (anonymous):

no it cant be factored

OpenStudy (solomonzelman):

yes it can be factored !

OpenStudy (anonymous):

ok..

OpenStudy (solomonzelman):

Think like this... you know that the first terms would be (x ) (x ) this way you will get x^2 in the beginning. you can also tell that because you have a negative middle term and a negative last term that one of the parenthesis will have a - in it, and the other will have a +. So (x - ) ( x + ) which 2 numbers would give us a Product (multiplication) of -10, and a sum of -3 ?

OpenStudy (solomonzelman):

\[10=(-5) \times 2=(-2) \times 5=(-10) \times 1=10 \times (-1)\] check these pairs of factors.

OpenStudy (anonymous):

nothing will give u that

OpenStudy (solomonzelman):

You are pushing me to do, alright... ! What about \[(x+2)(x-5)\] ?

OpenStudy (anonymous):

what about that

OpenStudy (solomonzelman):

Does that give you \[x^2-3x-10 ~~~~~~~~~~~~~~~?\]

OpenStudy (anonymous):

lol how on earth does make any sense

OpenStudy (anonymous):

no? yes?

OpenStudy (anonymous):

lol is this even what the question is asking

OpenStudy (solomonzelman):

I am just taking this as - you don't know how to factor. I would really suggest you to contact your teacher whether you are schooled online, or anywhere else. You got to know this staff, you won't be able to proceed in life without it. I would be willing to teach you, but you are not willing to get help or guidance, you just want to say this is the answer because, and you will take the answer without the because. I am really sorry to say all of this. I am also a teenager just like you, and believe me, I know how hard it is without a teacher.

OpenStudy (solomonzelman):

The question is asking you which option is correct. I was just helping you find the x intercepts, but you knowledge seems too small to do this.

OpenStudy (anonymous):

oh it is, havent you seen my username ? i might not know how to factor but i damn sure know how to do things that will make me successful, dont confuse your ability to factor with your pride honey, have a good day